# Towards Robust Zero-Shot Reinforcement Learning

### v1

@@ B.7 Hyperparameter
Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$
Walker-RND 9 0.1 64 0.99 0.001 5
Walker-APS 9 0.1 64 0.99 0.001 5
Walker-PROTO 9 0.1 64 0.99 0.001 5
Walker-DIAYN 9 0.1 64 0.99 0.001 5
Jaco-RND 1 0 64 0.99 0.001 10
Jaco-APS 9 0.1 64 0.99 0.0001 5
Jaco-PROTO 1 0 64 0.99 0.001 10
Jaco-DIAYN 9 0.1 64 0.99 0.0001 5
Quadruped-RND 1 0 64 0.99 0.001 10
Quadruped-APS 1 0 64 0.99 0.001 10
Quadruped-PROTO 1 0 64 0.99 0.001 10
Quadruped-DIAYN 1 0 64 0.99 0.001 10
Kitchen-mixed 1 0 16 0.7 0.001 5
Kitchen-partial 2 0.2 8 0.7 0.001 5
@@
Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$
Walker-RND 2 0.2 32 0.99 0.0001 10
Walker-APS 2 0.2 32 0.99 0.0001 10
Walker-PROTO 2 0.2 32 0.99 0.001 10
Walker-DIAYN 8 0.2 32 0.99 0.001 10
Jaco-RND 1 0 32 0.99 0.0001 10
Jaco-APS 1 0 32 0.99 0.0001 10
Jaco-PROTO 1 0 32 0.99 0.0001 10
Jaco-DIAYN 1 0 32 0.99 0.0001 10
Quadruped-RND 2 0.2 32 0.99 0.001 10
Quadruped-APS 1 0 32 0.99 0.001 10
Quadruped-PROTO 2 0.2 32 0.99 0.001 10
Quadruped-DIAYN 2 0.2 32 0.99 0.0001 10
@@ D.1 Learning Curves
This section presents the zero-shot learning curves on the ExORL benchmark. Results on the full dataset and the 100k small-sample dataset are shown in Figure 9-11 and Figure 12-14, respectively. The solid lines represent the mean IQM returns over 5 random seeds, and the shaded regions correspond to the standard deviation.
@@ D.2 Detailed Results on ExORL Benchmarks
This section presents complete experimental results on the ExORL benchmark. Table 8 summarizes performance on the full dataset, while Table 9 reports results on the 100k small-sample datasets. Evaluations were conducted every 10,000 steps during offline training. Following standard zero-shot RL criteria, we computed the aggregate score across all tasks as the overall performance measure, and selected the highest domain-level score across 5 random seeds along with corresponding per-task results.
@@ D.3 Empirically Evaluated $M^{\pi_{z}}$ and $Q_{z}$ on Other Tasks
We present empirical visualization results of $M^{\pi_{z}}$ and $Q_{z}$ distributions, shown in Figure 18-18. We trained FB, VCFB, and MCFB on the RND [5] dataset. Specifically, the trained $B$ network was utilized for inferring the task latent vector $z_{\text{task}}$, and as the input to both $M^{\pi_{z}}$ and $Q_{z}$ during the evaluation phase.

### v2

@@ B.7 Hyperparameter
Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$ Temperature $\alpha$
Walker-RND 9 0.1 64 0.99 0.001 5 0.05
Walker-APS 9 0.1 64 0.99 0.001 5 0.05
Walker-PROTO 9 0.1 64 0.99 0.001 5 0.05
Walker-DIAYN 9 0.1 64 0.99 0.001 5 0.05
Jaco-RND 1 0 64 0.99 0.001 10 0.05
Jaco-APS 9 0.1 64 0.99 0.0001 5 0.05
Jaco-PROTO 1 0 64 0.99 0.001 10 0.05
Jaco-DIAYN 9 0.1 64 0.99 0.0001 5 0.05
Quadruped-RND 1 0 64 0.99 0.001 10 0.05
Quadruped-APS 1 0 64 0.99 0.001 10 0.05
Quadruped-PROTO 1 0 64 0.99 0.001 10 0.05
Quadruped-DIAYN 1 0 64 0.99 0.001 10 0.05
Kitchen-mixed 1 0 16 0.7 0.001 5 0.1
Kitchen-partial 2 0.2 4 0.7 0.001 5 0.08
@@
Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$ Temperature $\alpha$
Walker-RND 2 0.2 32 0.99 0.0001 10 0.05
Walker-APS 2 0.2 32 0.99 0.0001 10 0.05
Walker-PROTO 2 0.2 32 0.99 0.001 10 0.05
Walker-DIAYN 8 0.2 32 0.99 0.001 10 0.05
Jaco-RND 1 0 32 0.99 0.0001 10 0.05
Jaco-APS 1 0 32 0.99 0.0001 10 0.05
Jaco-PROTO 1 0 32 0.99 0.0001 10 0.05
Jaco-DIAYN 1 0 32 0.99 0.0001 10 0.05
Quadruped-RND 2 0.2 32 0.99 0.001 10 0.05
Quadruped-APS 1 0 32 0.99 0.001 10 0.05
Quadruped-PROTO 2 0.2 32 0.99 0.001 10 0.05
Quadruped-DIAYN 2 0.2 32 0.99 0.0001 10 0.05
@@ D.1 Learning Curves
This section presents the zero-shot learning curves on the ExORL benchmark. Results on the full dataset and the 100k small-sample dataset are shown in Figure 10-12 and Figure 13-15, respectively. The solid lines represent the mean IQM returns over 5 random seeds, and the shaded regions correspond to the standard deviation.
@@ D.2 Detailed Results on ExORL Benchmarks
This section presents complete experimental results on the ExORL benchmark. Table 7 summarizes performance on the full dataset, while Table 8 reports results on the 100k small-sample datasets. Evaluations were conducted every 10,000 steps during offline training. We computed the aggregate score across all tasks as the overall performance measure, and selected the highest domain-level score across 5 random seeds along with corresponding per-task results.
@@ D.3 Empirically Evaluated $M^{\pi_{z}}$ and $Q_{z}$ on Other Tasks
We present empirical visualization results of $M^{\pi_{z}}$ and $Q_{z}$ distributions, shown in Figure 19-19. We trained FB, VCFB, and MCFB on the RND [5] dataset. Specifically, the trained $B$ network was utilized for inferring the task latent vector $z_{\text{task}}$, and as the input to both $M^{\pi_{z}}$ and $Q_{z}$ during the evaluation phase.